# Debt, Growth, and the Carbon Lock-In

### v1

@@ Results
Using strategy B, the solvency probability cannot be approximated as for strategy A. We obtain instead
@@
with $X_{t_{p}}=\Pi_{t_{p}}-(L-1)\sum_{t\leq t_{p}-1}\Pi_{t}/\rho$ where $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$. We observe in Fig.4D that it is a decreasing function of time, as shown in the SI. Interestingly, this result holds true whenever $\left<\gamma\right><\rho$ and for low enough leverages. Thus, we recover the inequality that sets the stability of public debt with respect to GDP [34]. From this, we conclude that debt tends to rise faster on average than capital as long as the intrinsic growth rate is smaller than the interest rate $\rho$.
Typically, solvency is lower under strategy B because a higher risk must be taken to sustain a higher growth rate. For $L\sim 1$ we understand why this solvency probability (strategy B) is smaller than for strategy A, as the lower bound for solvency is larger. We also show in the SI that increasing leverage (thus relying more on credit) can be considered ”optimal” because it increases the growth rate of $C_{\tau}$, whereas the survival probability plateaus at large leverage. However, relying on such a strategy implies stronger carbon emissions, particularly for carbon intensive economies, as discussed above.
Effect of leverage on solvency (strategy B)
The risk associated with growth stochasticity can be evaluated using the probability of solvency. In particular, we can examine the impact of leverage and payback time on the probability of solvency. On Fig.4 we see that solvency probability generally decreases with both leverage and payback time in the range of leverage corresponding to real data. This means that debt is used as a means to increase short-term GDP, even if it results in increased long-term risk, which is measured by cumulative carbon emissions (depending on the economy’s carbon intensity) and a debt that grows faster than GDP.
In SI 2.D and 2.E, we further explore solvency in both models, and compute the optimal payback time for the post-payback growth rate conditional on solvency.
Figure 4: Illustration of trade-off between solvency and payback time. Solvency probability $P_{S}$ is shown for model B with the color scale as function of payback time $t_{p}$ and interest rate $\rho$. Solvency becomes less likely with time in this model whenever $\left<\gamma\right><\rho$ for all leverages $L$, i.e. when debt increases faster than the capital. This means that this strategy is not viable on long timescales, and is just a way to increase short-term GDP (resulting in increased carbon emissions).
Discussion
Our analysis reveals that debt-financed growth acts as a structural amplifier of emissions: leveraging $L$ boosts GDP $C_{\tau}$ and, given the slow evolution of carbon intensity $I(\tau)$, results in cumulative emissions that remain nearly proportional to cumulative GDP, confirming previous studies [37][54][55]. By separating the growth due to deficits from the intrinsic growth, we have confirmed that the economic growth of several countries is closely tied to deficits.
We calibrated the model over multi-decade windows using observed macroeconomic data, extracting each country’s effective leverage $L$ (the average ratio of deficit-financed spending to GDP), the intrinsic growth rate $\gamma$ (the underlying growth rate absent new borrowing), and an interest rate $\rho$. This data-driven approach is not a free fit, but rather implements recorded deficits, public debt, and GDP to set model parameters. Therefore, the model’s predictions can be directly compared to the actual outcomes.
To our knowledge, this is the first model to explicitly link fiscal leverage with cumulative CO2 emissions and to calibrate it using multi-decade country-level data. Indeed, economic growth is typically treated as exogenous or as the result of a production function in Integrated Assessment Models (IAMs), such as DICE, REMIND, or MESSAGEix, without explicitly modelling credit dynamics [56, 57]. Similarly, the majority of climate-finance studies focus on transition risks or stranded assets [17], rather than quantifying the structural amplification of GDP growth and emissions by leverage. Our method is similar to macro-energy analyses that connect debt and energy consumption [2, 58]. However, it goes further by calibrating leverage $L$, intrinsic growth $\gamma$, and interest rate $\rho$ based on historical data. This allows for the direct prediction and validation of cumulative emissions trajectories. For all four considered countries, the model accurately tracks the cumulative CO2 emissions per capita alongside economic growth (Fig. 2B and SI), confirming that cumulative emissions are closely coupled with cumulative GDP, thus resulting in a persistent carbon lock-in effect. We can observe that, despite differences in decarbonization efforts, each additional unit of GDP has historically carried a largely invariant carbon cost over the long run (within the calibration period), carrying near-proportional relationships between total economic output and total emissions. For example, using a time-declining carbon intensity $I(t)$ for the US, the integration of our model from 1990 to 2020 yields cumulative per capita CO2 emissions of the order $7\times 10^{5}$ kg, in good agreement with the observed value of $\sim 6\times 10^{5}$ kg. A similar close agreement is found for China, France, and Denmark (Fig. 2B), as evidenced by the overlapping model trajectories and historical data points. This level of precision indicates that our leveraged growth model captures a first-order causal driver of emissions: debt-fuelled economic expansion. Hence, it can provide a partial explanation for why emissions have continued to increase in line with commodity production, despite gains in energy efficiency in many sectors. The differences between countries in economic structure and debt dynamics are reflected in the calibrated parameters and outcomes. We find that the United States is heavily dependent on deficit spending to sustain its growth. Indeed, since 1980, its intrinsic growth rate has averaged to only one percent per year, with substantial variability. In other words, most of the US GDP growth since 1980 can be attributed to credit expansion (public deficit and debt accrual) rather than to productivity gains [59]. Our calibration leads to an average leverage slightly above 1 for the US, rising to $L\approx 1.06$ in 2023 amid large federal spending. This credit-driven growth is reflected in the model’s ability in reproducing the persistent rise in cumulative U.S. CO2 emissions at the end of the 90s until just before the global financial crisis of 2008 (GFC) [60].
An additional insight from our framework is that debt-driven growth imposes financial risks and constraints. Our stochastic analysis of the solvency probability (Fig. 4 and SI) shows that when the intrinsic growth rate $\langle\gamma\rangle$ falls below the interest rate $\rho$, the probability of long-term solvency decays towards zero, regardless of leverage $L$. In other words, maintaining growth through excessive borrowing ultimately becomes unsustainable and carries an endogenous risk of default. This result provides a natural constraint on fiscal stimulus strategies: even before critical climate transitions are reached, macro-financial dynamics can force deleveraging and major systemic destabilisations. From a control theory perspective, this identifies an optimal leverage level on the Pareto front, beyond which additional credit raises default risk more than it supports sustainable growth [62, 63, 23].
One key limitation of the model is its tight coupling of public debt and GDP. We assume that each country’s government borrows a fixed fraction of GDP each year (the leverage $L$ above 1 corresponds to a primary deficit ratio) and that this borrowing directly and immediately adds to that year’s GDP. In reality, the relationship between debt and growth is more complex: private credit creation (by banks and firms) can also drive variability, and not all deficit spending translates into GDP in the short run (in fact, the fiscal multiplier can vary). Our model effectively treats deficit-financed spending as equally productive as other spending, whereas inefficient investments or financial sector leakages could reduce the growth payoff of debt. In addition, identifying the primary deficit with new capital investment is an oversimplification. In fact, governments may borrow for many reasons, and not all borrowing fuels productive capacity. A more advanced model would require integrating both public and private debt dynamics and the ability to distinguish between productive investment and pure demand stimulus.
Denmark presents another informative and distinctive boundary case for our framework, owing to its limited public borrowing and persistent fiscal surpluses. As illustrated in the SI, Denmark’s territorial emissions have decreased, while its consumption-based emissions have remained relatively constant, indicating that no actual decoupling has occurred. This discrepancy suggests that fiscal restraint alone is insufficient to guarantee a reduced carbon footprint when emissions are outsourced through trade. In the SI, we compare territorial and consumption-based estimates (Fig. S4) and find that offshoring of emissions can significantly alter apparent decoupling trends (a phenomenon also referred to as carbon leakage [69]). Despite their high consumption-based footprint, countries that relocate carbon-intensive production abroad appear to decrease their emissions-to-GDP ratio [70]. In a similar spirit, network-based allocation methods developed by Van den Ende et al. (2025) offer a rigorous approach to quantify responsibility for emissions along global value chains, which complements territorial accounting [71].
Financial debt and carbon debt are two sides of the same anticipatory system: the former brings future wealth forward, while the latter consumes the energy of the future. Kelly’s model, when scaled up to the macroeconomic level, shows that there are both physical and financial limits to debt-driven growth. It is an important message because the idea that there are limits to growth, even when allowing for innovation, is not widely accepted among economists [75]. Here, we show that as long as credit rests on the promise of material expansion, each euro borrowed creates a double debt — a financial and a climatic one. The great challenge of the twenty-first century will be to decouple prosperity from the credit lever, so that the growth of well-being is no longer tied to the growth of carbon.
Methods
We report the parameters used in the model to simulate the evolution of the $GDP$ in Table 1. We averaged their values over the interval 199820221998-2022 to use as inputs in our model.
@@ Methods
$\left<L_{\tau}\right$> $1.025\pm 0.023$ $1.051\pm 0.040$ $0.997\pm 0.029$ $1.040\pm 0.018$
$\left<\rho_{\tau}\right$> $1.131\pm 0.101$ $1.013\pm 0.051$ $1.018\pm 0.237$ $1.008\pm 0.088$
$W$ $088\pm 0.086 0.0$$18\pm 0.059 0.0$$32\pm 0.089 0.0$$020\pm 0.090$
$\left<\gamma_{\tau}\right$> $1.096\pm 0.095$ $0.983\pm 0.056$ $1.036\pm 0.091$ $0.983\pm 0.088$
Table 1: Values of the different parameters.
We also report the coefficients of determination $R^{2}$ and rescaled Mean Absolute Errors obtained between data and our model for Fig.2 A and B in table.2 and table.3. The model is used to simulate the evolution of the GDP with its related stochasticity, and uses measured yearly carbon intensities as an input to simulate path-dependent intensities $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$.
China US Denmark France
$R^{2}_{\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)}$ $0.985$ $0.938$ $0.981$ $0.982$
$R^{2}_{\mathcal{E}(\tau)}$ $0.980$ $0.996$ $0.982$ $0.977$
Table 2: Coefficients of determination $R^{2}$ between data and theory for path-dependent intensities and cumulative emissions per capita, for different countries. With $(X_{n})$ being the real data and $X_{n}^{\text{simu}}$ the simulated data, we have $R^{2}_{X}=1-\sum\sum_{n=1}^{N_{m}}\left((X_{n}-X_{n}^{\text{simu}})/(X_{n}-\left<X\right>)\right)^{2}$.
China US Denmark France
$MAE_{\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)}/\left<\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right$> $0.032$ $0.035$ $0.027$ $0.017$
$MAE_{\mathcal{E}(\tau)}/\left<\mathcal{E}(\tau)\right$> $0.092$ $0.027$ $0.057$ $0.057$
Table 3: Rescaled Mean Absolute Error MAEMAE between data and theory for path-dependent intensities and cumulative emissions per capita, for different countries. This quantity is computed as $MAE_{X}/\left<X\right>=\left(\sum_{n=1}^{N_{m}}|X_{n}-X_{n}^{\text{simu}}|\right)/\left(\sum_{n=1}^{N_{m}}X_{n}\right)$ with $N_{m}$ the number of measurements.

### v2

@@ Results
In the case of strategy B, unlike the case of strategy A, the solvency probability cannot be written in a closed form. We derive instead a recursive equation for the net capital from which we deduce that this solvency probability is a decreasing function of leverage provided that the intrinsic growth rate is lower than the interest rate and leverage is sufficiently close to $1$. We observe this trend in Fig. 4 and provide details of the calculations in Methods section and in the Supplementary Information.
Typically, solvency is lower under strategy B because a higher risk must be taken to sustain a higher growth rate. For $L_{\tau}\sim 1$, this solvency probability (strategy B) is smaller than for strategy A, since the investor relies more on leverage in strategy B than in A.
We then examine the impact of leverage and payback time on the probability of solvency. As expected, we observe that solvency probability generally decreases with both leverage and payback time, as illustrated in Fig. 4. In other words, debt leads to short-term GDP gains, even if it results in an increased long-term risk and enhanced cumulative carbon emissions, especially when the debt grows faster than GDP. In SI 2.4 and 2.5, we further explore solvency in both models. Notably, we find that high leverage strategies are optimal when the objective is the maximization of the growth rate for a given level of risk as measured by the solvency probability. However, these strategies are only optimal because environmental costs are ignored. In practice, high leverage strategies typically lead to large carbon emissions, particularly for carbon intensive economies, as discussed above.
Figure 4: Illustration of trade-off between solvency and payback time. Solvency probability $P_{S}$ is shown for model B with the color scale as function of payback time $t_{p}$ and interest rate $\rho-1$. Solvency becomes less likely with time in this model whenever $\left<\gamma\right><\rho$ for all leverages $L$, i.e. when debt increases faster than the capital. This means that this strategy is not viable on long timescales, and is just a way to increase short-term GDP (resulting in increased carbon emissions).
Discussion
Our analysis reveals that debt-financed growth structurally amplifies emissions. Specifically, the leverage $L$ boosts GDP $C_{\tau}$, while the slow decline in carbon intensity $I(\tau)$ leads to cumulative emissions that remain nearly proportional to cumulative GDP. This confirms previous studies [41][59][60]. By separating the growth due to deficits from the intrinsic growth, we have confirmed that the economic growth of several countries is closely tied to deficits.
We calibrated the model over multi-decade windows using observed macroeconomic data, extracting each country’s effective leverage $L$ (the average ratio of deficit-financed spending to GDP), the intrinsic growth rate $\gamma$ (the underlying growth rate absent new borrowing), and an interest rate $\rho-1$. This data-driven approach is not a free fit, but rather implements recorded deficits, public debt, and GDP to set model parameters. Therefore, the model’s predictions can be directly compared to the actual outcomes.
To our knowledge, this is the first model to explicitly link fiscal leverage with cumulative CO2 emissions and to calibrate it using multi-decade country-level data. Indeed, economic growth is typically treated as exogenous or as the result of a production function in Integrated Assessment Models (IAMs), such as DICE, REMIND, or MESSAGEix, without explicitly modelling credit dynamics [61, 62]. Similarly, the majority of climate-finance studies focus on transition risks or stranded assets [20], rather than quantifying the structural amplification of GDP growth and emissions by leverage. Our method is similar to macro-energy analyses that connect debt and energy consumption [2, 63]. However, it goes further by calibrating leverage $L$, intrinsic growth $\gamma$, and interest rate $\rho-1$ based on historical data. This allows for the direct prediction and validation of cumulative emissions trajectories. For all four considered countries, the model accurately tracks the cumulative CO2 emissions per capita alongside economic growth (Fig. 2B and SI), confirming that cumulative emissions are closely coupled with cumulative GDP, thus resulting in a persistent carbon lock-in effect. We can observe that, despite differences in decarbonization efforts, each additional unit of GDP has historically carried a largely invariant carbon cost over the long run (within the calibration period), carrying near-proportional relationships between total economic output and total emissions. For example, using a time-declining carbon intensity $I(t)$ for the US, the integration of our model from 1990 to 2020 yields cumulative per capita CO2 emissions of the order $7\times 10^{5}$ kg, in good agreement with the observed value of $\sim 6\times 10^{5}$ kg. A similar close agreement is found for China, France, and Denmark (Fig. 2B), as evidenced by the overlapping model trajectories and historical data points. This level of precision indicates that our leveraged growth model captures a first-order causal driver of emissions: debt-fuelled economic expansion. Hence, it can provide a partial explanation for why emissions have continued to increase in line with commodity production, despite gains in energy efficiency in many sectors. The differences between countries in economic structure and debt dynamics are reflected in the calibrated parameters and outcomes. We find that the United States is heavily dependent on deficit spending to sustain its growth. Indeed, since 1980, its intrinsic growth rate has averaged to only one percent per year, with substantial variability. In other words, most of the US GDP growth since 1980 can be attributed to credit expansion (public deficit and debt accrual) rather than to productivity gains [64]. Our calibration leads to an average leverage slightly above 1 for the US, rising to $L\approx 1.06$ in 2023 amid large federal spending. This credit-driven growth is reflected in the model’s ability in reproducing the persistent rise in cumulative US CO2 emissions at the end of the 90s until just before the global financial crisis of 2008 (GFC) [65].
An additional insight from our framework is that debt-driven growth imposes financial risks and constraints. Our stochastic analysis of the solvency probability (Fig. 4 and SI) shows that when the intrinsic growth rate $\langle\gamma\rangle$ falls below the interest rate $\rho-1$, the probability of long-term solvency decays towards zero, regardless of leverage $L$. In other words, maintaining growth through excessive borrowing ultimately becomes unsustainable and carries an endogenous risk of default. This result imposes a natural constraint on fiscal stimulus strategies. Even before critical climate transitions are reached, macro-financial dynamics may trigger deleveraging and systemic destabilizations. From a control theory perspective, this identifies an optimal leverage level on the Pareto front, beyond which additional credit raises default risk more than it supports sustainable growth [67, 68, 28].
One key limitation of the model is its tight coupling of public debt and GDP. We assume that each country’s government borrows a fixed fraction of GDP each year (the leverage $L$ above 1 corresponds to a primary deficit ratio) and that this borrowing directly and immediately adds to that year’s GDP. In reality, the relationship between debt and growth is more complex: private credit creation (by banks and firms) can also drive variability, and not all deficit spending translates into GDP in the short run (in fact, the fiscal multiplier can vary). Our model effectively treats deficit-financed spending as equally productive as other spending, whereas inefficient investments or financial sector leakages could reduce the growth payoff of debt. In addition, identifying the primary deficit with new capital investment is an oversimplification. In fact, governments may borrow for many reasons, and not all borrowing fuels productive capacity. A more advanced model would require integrating both public and private debt dynamics and the ability to distinguish between productive investment and pure demand stimulus. Such a complementary analysis of private debt could be conducted, as done previously with macroeconomic models of financial instability [3, 7].
Denmark presents another informative and distinctive boundary case for our framework, owing to its limited public borrowing and persistent fiscal surpluses. As illustrated in the Supplementary Information, Denmark’s territorial emissions have decreased, while its consumption-based emissions have remained relatively constant, indicating that no actual decoupling has occurred. This discrepancy suggests that fiscal restraint alone is insufficient to guarantee a reduced carbon footprint when emissions are outsourced through trade. In the Supplementary Information, we compare territorial and consumption-based estimates (Fig. S2) and find that offshoring of emissions can significantly alter apparent decoupling trends (a phenomenon also referred to as carbon leakage [74]). Despite their high consumption-based footprint, countries that relocate carbon-intensive production abroad appear to decrease their emissions-to-GDP ratio [75]. To address this issue, network-based allocation methods developed by Van den Ende et al. (2025) offer a rigorous approach to quantify responsibility for emissions along global value chains, which complements territorial accounting [76].
Financial debt and carbon debt are two sides of the same anticipatory system: the former brings future wealth forward, while the latter consumes the energy of the future. Our stochastic credit-risk model, when scaled up to the macroeconomic level, shows that there are both physical and financial limits to debt-driven growth. It is an important message because the idea that there are limits to growth, even when allowing for innovation, is not widely accepted among economists [80]. Here, we show that as long as credit rests on the promise of material expansion, each euro borrowed creates a double debt — a financial and a climatic one. The great challenge of the twenty-first century will be to decouple prosperity from the credit lever, so that the growth of well-being is no longer tied to the growth of carbon.
Methods
Definition of the stochastic credit-risk model
Our stochastic credit-risk model is defined by the following dynamical equations for the capital $C_{\tau}$ and the debt $D_{\tau}$ at time $\tau$:
$C_{0}$ $=B_{0}+C_{\text{ini}}=LC_{\text{ini}},$ (6)
$C_{\tau+1}$ $=\gamma_{\tau}C_{\tau}+B_{\tau+1}-\left(\sum_{\tau^{\prime}=0}^{\tau}\delta_{\tau,t_{p}(\tau^{\prime})}\right)D_{\tau+1},$ (7)
$D_{0}$ $=B_{0}=(L-1)C_{\text{ini}},$ (8)
where in the first equation, $C_{\text{ini}}$ is the amount invested initially by the investor, and $B_{0}$ is the initial amount borrowed. The leverage is $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. The second equation describes the growth of the capital from time $\tau$ to time $\tau+1$. The model uses discrete time, as in Kelly’s original model. However, extending it to continuous time is straightforward. At each time step, the capital is multiplied by $\gamma_{\tau}$, which accounts for inflation [38]. This parameter represents the stochastic instantaneous growth rate at time $\tau$. The evolution of the debt is described by the last two equations. The interest rate $\rho_{\tau}-1$ is the factor by which the amount borrowed is multiplied at each time step, in other words, the debt due to the amount borrowed at time $\tau^{\prime}$, is multiplied by $\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}$ if it has not been paid back at time $\tau$. In general, the interest rate is larger than 0 ($\rho>1$), so that the debt keeps increasing, and the borrower ends up paying more than the amount borrowed. If $\gamma_{\tau}>1$, the gross capital increases provided that there is no borrowing or payback at this time. The second term represents the amount borrowed at time $\tau$; it can be 0 if the borrower does not borrow at this step, or a positive value. The third term represents the possible payback of the debt (debt clearing) at time $\tau+1$, and will either be equal to 0 (no payback), or to $D_{\tau+1}$.
In the second and fourth equation, $t_{p}(\tau^{\prime})$ indicates the time when the amount borrowed at $\tau^{\prime}$ is paid back. In the second equation, if a payback happens at time $\tau$, $\exists\tau^{\prime},\tau=t_{p}(\tau^{\prime})$, and thus $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=1$, then the capital is reduced by the service of the debt, otherwise $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=0$ and there is no such reduction. In the last equation, the Heaviside function $\Theta$ function allows us to know if the payback has occurred at time $\tau$ (indeed $\Theta(t_{p}(\tau^{\prime})-\tau)=0$ if $\tau>t_{p}$, meaning that the debt is erased after it has been payed back).
Strategy A
We assume that the investor initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. We refer to this strategy as strategy A, which is akin to an individual taking out a loan to buy a house. In practice, the investor can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and the leverage is then $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho-1$. At the time of payback, the capital can become negative, which could not happen within the purely multiplicative dynamics of Kelly’s model. Bankruptcy occurs if $C_{t_{p}}$ is negative, or equivalently if
$\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}<1-\frac{1}{L},$ (10)
meaning that the larger $L$ is, the more likely the bankruptcy. Under the assumption of no bankruptcy, we can also define the long-term growth rate $W=\mathbb{E}\left({\ln\left(\gamma\right)}\right)$. Note that due to the concavity of the logarithm, we have generally for any probability distribution of $\gamma$, $\ln(E)\geq W$, with $E=\mathbb{E}\left({\gamma}\right)$. However, there is a non-zero probability of hitting bankruptcy when the debt is paid back (at $t_{p}$).
Strategy B
Within strategy B, the investor borrows to an external actor at each time step with a time dependent leverage $L_{\tau}$. The dynamics is then (except for the time of payback):
Now, in addition to the long-term growth rate $W$ defined as before, one can also consider the short-term growth rate $\omega(t_{p})$ (defined in SI section 1.1). We explicitly separate the value linked to production from the value added through borrowing, in the stochastic intrinsic growth rate:
Solvency in Strategy A
The solvency probability $P_{S}(t_{p})$ for a fixed payback time $t_{p}$ can be computed explicitly in this case. To do so, we introduce the stochastic variable $X_{\tau^{\prime}}=\ln(\gamma_{\tau^{\prime}}/\rho)$, which represents increments of the long-term growth rate. This variable has mean $m=W-\ln(\rho)$ and variance $\sigma_{X}^{2}$. Then, we show in Supplementary Information that
$P_{S}(t_{p})=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right).$ (13)
From this equation, we find that three regimes are possible depending on the sign of $m$. If $m>0$, $P_{S}(t_{p}\to\infty)\to 1$, solvency is guaranteed, and there is thus little pressure to pay back the debt. If $m<0$ instead, $P_{S}(t_{p}\to\infty)\to 0$ and the convergence is faster as $\sigma_{X}$ decreases, but in this case, the best course of action is to pay back the debt as soon as possible to avoid bankruptcy. The case $m=0$ is a limiting case where $P_{S}(t_{p}\to\infty)=\frac{1}{2}$.
Solvency in Strategy B
The expression for the solvency probability in this case is
@@ Methods
with $X_{t_{p}}$ now defined by $X_{t_{p}}=\Pi_{t_{p}}-(L-1)\sum_{t\leq t_{p}-1}\Pi_{t}/\rho$ with $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$. We identify two regimes in this case, which are controlled by the sign of $\left<\gamma\right>-\rho$. Under the assumption that fluctuations in the economy are not too large, we expect the solvency probability to be a decreasing function of time when $\left<\gamma\right><\rho$, whereas the opposite is expected for $\left<\gamma\right>>\rho$. Thus, we recover a well-known condition that sets the stability of public debt with respect to GDP [38]. From this, we conclude that debt tends to rise faster on average than capital as long as the intrinsic growth rate is smaller than the interest rate $\rho-1$.
Data analysis
We have extracted from economic data of four countries (China, US, Denmark and France) in the time frame 199820221998-2022, parameters of the model as shown in Table 1.
@@
Average leverage $\left<L_{\tau}\right$> $1.025\pm 0.023$ $1.051\pm 0.040$ $0.997\pm 0.029$ $1.040\pm 0.018$
Average interest rate $\left<\rho_{\tau}\right>-1$ $131\pm 0.101 0.0$$13\pm 0.051 0.0$$18\pm 0.237 0.0$$008\pm 0.088$
Average long-term growth rate $W$ $088\pm 0.086 0.0$$18\pm 0.059 0.0$$32\pm 0.089 0.0$$020\pm 0.090$
Average intrinsic growth rate $\left<\gamma_{\tau}\right>-1$ $0.096\pm 0.095$ $17\pm 0.056 0.0$$36\pm 0.091 0.0$$017\pm 0.088$
Table 1: Estimations of key parameters of the model from economic data of GDP and deficits of four different economies.